OpenStudy (anonymous):

Plastic material worth 2.75\kg is to be mixed with plastic worth 3.50\kg to make up a 50 kg mixture to sell at 3 kg.

OpenStudy (anonymous):

Let the amount of 1st plastic needed = x kg Let the amount of 2nd plastic needed = y kg therefore by data x+y=50 -1st eq Therefore also by data, 2.75x+3.5y=3(50) 2.75x +3.5y=150 -2nd eq Multiply the second equation by 100 275x+350y=15000 Now simplifying this equation we have, 11x + 14y =600 -3rd eq Multiply the first equation by 11, 11x +11y=550 -4th Therefore subtracting the 4th eq from the 3rd we have, 3y=50 y=50/3=16.67kg therefore x=150-50/3=100/3=33.33kg Hence you need to mix 16.67 kg of the first plastic and 33.33kg of the second plastic !!

OpenStudy (anonymous):

3.5k + 2.75 ( 50 -k) = 3 * 50 -> .75k = 150 - 137.5 => K = 12.5/ .75 = 16.7 kg of $3.5/kg

OpenStudy (anonymous):

and 33.3 kgs of $2.5/kg